### v1

@@ 5.1 Sequence-to-Sequence Language Translation
We consider a sequence-to-sequence language translation task on the WMT16 Romanian-English dataset Bojar et al. (2016), using BLEU Papineni et al. (2002) and ROUGE-L Lin (2004) as the objectives. Following Peters & Martins (2021), the dataset is preprocessed with SentencePiece tokenization Kudo (2018), and an LSTM-based encoder-decoder is trained. Two key hyperparameters are considered:
1. The hyperparameter $\rho$ controls the sparsity of the output distribution using Entmax Peters et al. (2019), transitioning between a dense output with softmax ($\rho=1$) and sparsemax ($\rho=2$) Martins & Astudillo (2016).
2. The Fenchel-Young label smoothing strength $\epsilon$ is a training regularization hyperparameter that determines the extent to which one-hot targets are mixed with uniform noise based on Fenchel-Young losses Peters & Martins (2021). Accordingly, the original one-hot targets are assigned weight $1-\epsilon$, while the uniform distribution over all possible classes is assigned the weight $\epsilon$.
To create the initial candidate set $\Lambda$, we selected hyperparameters over a grid of 32 combinations, using 8 logarithmically spaced values in the interval $[1,2]$ for $\rho$, and values in the set $\{0.0,0.01,0.05,0.1\}$ for $\epsilon$. This selection is in line with Peters & Martins (2021).
@@
5.2 Image Segmentation for Object Detection
We now evaluate the proposed RG-PT framework on a multi-objective image segmentation task for object detection, leveraging the MS-COCO dataset Lin et al. (2014) and a pretrained detector from Detectron2 Wu et al. (2019) as done in Angelopoulos et al. (2021). The task involves three distinct objectives: (i) detecting objects within an image (object detection); (ii) delineating object boundaries (image segmentation); and (iii) assigning correct labels to detected objects (object classification). These tasks are measured using recall, intersection-over-union (IoU), and classification accuracy, respectively. The goal is to control classification errors while optimizing recall and segmentation quality, addressing the trade-offs among these objectives.
Denote as $R_{1}(\lambda)$, $R_{2}(\lambda)$, and $R_{3}(\lambda)$ the risks associated with recall, IOU, and coverage, respectively, for hyperparameter $\lambda=(\lambda_{1},\lambda_{2},\lambda_{3})$. Controlling these risks in the context of problem (3) is equivalent to having $L=3$ and $L_{c}=3$. Additionally, as in Angelopoulos et al. (2021), we set the targets as $\alpha_{1}=0.5$, $\alpha_{2}=0.5$, and $\alpha_{3}=0.75$ with $\delta=0.1$. Within the set $\hat{\Lambda}_{\mathcal{Z}}$ of reliable hyperparameters returned by the algorithm of choice, we choose the hyperparameter in subset $\hat{\Lambda}_{\mathcal{Z}}$ with the lowest value of $\lambda_{1}$ in order to increase the number of detected objects as much as possible. No prior knowledge was leveraged in creating the RG in this experiment by setting $n_{p}=0$.
6 Conclusion
@@
Future work may investigate the use of synthetic data for the derivation of an RG, as well as the integration with sequential testing methods based on e-processes Zecchin & Simeone (2024).
Acknowledgements
@@ Appendix A Fixed Sequence Testing
In this section, we provide a brief overview of FST for controlling the FDR, which we used in our simulations for PT. While PT, as outlined in Laufer-Goldshtein et al. (2023), is designed to support control of the FWER, our focus in this paper is on FDR control.
MHT methods such as the Bonferroni correction for FWER control Rice (2006) and the Benjamini-Yekutieli (BY) procedure for FDR control Benjamini & Yekutieli (2001) do not leverage any side information about the relative reliability of the hyperparameters. When such information is available during calibration, FST can be used to test hyperparameters in order of expected reliability. When the ordering information is accurate, FST can be beneficial to reduce the FDR Lynch et al. (2017). In this section, we briefly describe the FST procedure for FDR control.
@@
The critical values $\delta_{i}$ are adapted to account for the position of each hypothesis in the testing sequence. These values are specifically designed to control the FDR under various dependency structures among the $p$-values. For the case of interest here, which is arbitrary dependence of the $p$-values, the critical levels can be set as Lynch et al. (2017)
@@ Appendix B Summary of DAGGER
This section outlines the step-up procedure used in DAGGER Ramdas et al. (2019) to dynamically adjust the testing thresholds. DAGGER determines the testing thresholds adaptively based on the structure of the DAG and on the outcomes of previously tested hypotheses. At each depth level of the DAG, thresholds are updated dynamically to control the FDR, while respecting the hierarchical dependencies encoded by the DAG.
@@
where $r\in[1,|\Lambda_{d}|]$ is a parameter set as detailed below; $v_{i}$ is the effective number of leaves in the subgraph rooted at the current node; $V$ is the total number of leaves in the DAG; $m_{i}$ is the effective number of nodes in the subgraph rooted at the current node; $R_{1:d-1}$ is the total number of rejections at depths $1$ through $d-1$; and $\beta(\cdot)$ is a reshaping function, such as the Benjamini-Yekutieli Benjamini & Yekutieli (2001) function $\beta_{BY}(x)=x\sum_{k=1}^{V}1/k$, which is designed to ensure FDR control under arbitrary dependence.
@@
Appendix C Additional Experiments
C.1 Image Classification
@@
In this experiment, following Franceschi et al. (2024), we consider the problem of hyperparameter selection for a support vector machine (SVM) model used to classify images from the Fashion MNIST dataset. The Fashion MNIST is a widely used benchmark for image classification, consisting of 70,000 grayscale images of 10 different clothing categories.
@@
This is a non-trivial problem since the accuracy maximizing model may not also optimize the recall Powers (2011).
The SVM model requires the selection of two hyperparameters Franceschi et al. (2024). The regularization parameter, $C$, controls the desired trade-off between maximizing the margin and minimizing the classification error. Lower values of $C$ allow for a softer margin that can overlook some misclassification errors, while higher values enforce stricter classification error requirements. The kernel coefficient, $\gamma$, determines the impact of a single training example on the decision boundary, with higher values capturing finer details in the data set but risking overfitting. To create the initial candidate set $\Lambda$, we selected hyperparameters over a grid of 25 combinations, using five logarithmically spaced values in the intervals $[-3,3]$ and $[-4,1]$ for $C$ and $\gamma$, respectively. This selection is in line with approaches such as Yogatama et al. (2015) for SVM hyperparameter selection.
@@
Fig. 8 illustrates the testing procedures of LTT Angelopoulos et al. (2021), PT Laufer-Goldshtein et al. (2023), and RG-PT. The x- and y-axes represent the logarithmic scales of the two hyperparameters $C$ and $\gamma$, while the contours indicate levels of recall $R_{\text{rec}}(\lambda)$, evaluated on the test data set, as a function of the hyperparameters $\lambda=(c,\gamma)$. The numbers illustrate the testing order for each testing method. Note that than LTT, which uses BY, does not follow any inferred order on the hyperparameters, and thus does not have the order labels in the figures. Furthermore, while LTT and PT test hyperparameters one by one, following a linear trajectory, RG-PT proceeds along a DAG, testing at the same time all hyperparameters at the same depth in the DAG.
@@
Figure 8: Illustration of the hyperparameter selection procedure followed by LTT (a), PT (b), and RG-PT (c) for the setting studied in Sec. C.1. Each node represents a hyperparameter $\lambda=(C,\gamma)$, with the numbers representing the testing order. Green nodes show the hyperparameters included in the reliable set $\hat{\Lambda}_{\mathcal{Z}}$, and the star node shows the hyperparameter in set $\hat{\Lambda}_{\mathcal{Z}}$ with the lowest recall rate.
C.2 Radio Access Scheduling
In this section, we study a telecommunications engineering problem, namely the optimization of a radio access scheduler Valcarce (2020). In this setup, each user equipment (UE) belongs to one of four quality-of-service (QoS) classes, assigned at random, each with its own delay and bit rate requirements de Sant Ana & Marchenko (2020). The goal is to control the delay of UEs in a given QoS class, while simultaneously minimizing the delays for UEs in the other three QoS classes.
@@ C.2 Radio Access Scheduling
The scheduling algorithm at the base station allocates spectral resources to the UEs. As in Valcarce (2020), the UEs are randomly distributed within a $1\,\text{km}^{2}$ area containing a centrally located base station. Each UE has an initial buffer of 100 packets, and moves at random speeds and directions. Resource allocation is carried out in intervals of 1 ms, called transmission time intervals (TTIs), over 10,000 TTIs per episode.
The resource allocation algorithm is controlled by a set of hyperparameters $\lambda=(\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4})\in\Lambda$, where each $\lambda_{i}$ adjusts a specific criterion in the reward model as detailed in Valcarce (2020). Hyperparameters $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$, and $\lambda_{4}$ determine respectively the channel quality for each UE, the total queue sizes at the UEs, the age of the oldest packet in each UE’s buffer, and the fairness in resource block allocation among UEs.
Calibration and test data were generated using the Nokia wireless suite Valcarce (2020). For each run, we used 100 episodes for calibration and 100 episodes for testing.
The candidate hyperparameter set $\Lambda$ was generated by keeping $\lambda_{1}$ and $\lambda_{2}$ at the values $\lambda_{1}^{*}$ and $\lambda_{2}^{*}$ recommended by de Sant Ana & Marchenko (2020), and linearly sweeping hyperparameters $\lambda_{3}$ and $\lambda_{4}$ in [0.02, 0.2] and [-0.1, 0.1], respectively, with 10 steps each, resulting in a total of 100 combinations.

### v2

@@ 4.2 Sequence-to-Sequence Language Translation
We consider a sequence-to-sequence language translation task on the WMT16 Romanian-English dataset [37], using BLEU [38] and ROUGE-L [39] as the objectives. Following [27], the dataset is preprocessed with SentencePiece tokenization [40], and an LSTM-based encoder-decoder is trained. Two key hyperparameters are considered:
1. The hyperparameter $\rho$ controls the sparsity of the output distribution using Entmax [41], transitioning between a dense output with softmax ($\rho=1$) and sparsemax ($\rho=2$) [42].
2. The Fenchel-Young label smoothing strength $\epsilon$ is a training regularization hyperparameter that determines the extent to which one-hot targets are mixed with uniform noise based on Fenchel-Young losses [27]. Accordingly, the original one-hot targets are assigned weight $1-\epsilon$, while the uniform distribution over all possible classes is assigned the weight $\epsilon$.
To create the initial candidate set $\Lambda$, we selected hyperparameters over a grid of 32 combinations, using 8 logarithmically spaced values in the interval $[1,2]$ for $\rho$, and values in the set $\{0.0,0.01,0.05,0.1\}$ for $\epsilon$. This selection is in line with reference [27].
@@
5 Conclusion, Limitations, and Future Work
@@
Limitations of this work include the exclusive applicability to settings with discrete hyperparameter spaces and the lack of theoretical results on the power and sample efficiency of the method. Future work may focus on optimizing the RG structure to maximize power, on the use of synthetic data for the derivation of an RG, as well as on the integration with sequential testing methods based on e-processes [43].
Acknowledgments
@@ Appendix A Fixed Sequence Testing
In this section, we provide a brief overview of FST for controlling the FDR, which we used in our simulations for PT. While PT, as outlined in [5], is designed to support control of the FWER, our focus in this paper is on FDR control.
MHT methods such as the Bonferroni correction for FWER control [23] and the Benjamini-Yekutieli (BY) procedure for FDR control [44] do not leverage any side information about the relative reliability of the hyperparameters. When such information is available during calibration, FST can be used to test hyperparameters in order of expected reliability. When the ordering information is accurate, FST can be beneficial to reduce the FDR [30]. In this section, we briefly describe the FST procedure for FDR control.
@@
The critical values $\delta_{i}$ are adapted to account for the position of each hypothesis in the testing sequence. These values are specifically designed to control the FDR under various dependency structures among the $p$-values. For the case of interest here, which is arbitrary dependence of the p-values, the critical levels can be set as [30]
@@ Appendix B Summary of DAGGER
This section outlines the step-up procedure used in DAGGER [15] to dynamically adjust the testing thresholds. DAGGER determines the testing thresholds adaptively based on the structure of the DAG and on the outcomes of previously tested hypotheses. At each depth level of the DAG, thresholds are updated dynamically to control the FDR, while respecting the hierarchical dependencies encoded by the DAG.
@@
where $r\in[1,|\Lambda_{d}|]$ is a parameter set as detailed below; $v_{i}$ is the effective number of leaves in the subgraph rooted at the current node; $V$ is the total number of leaves in the DAG; $m_{i}$ is the effective number of nodes in the subgraph rooted at the current node; $R_{1:d-1}$ is the total number of rejections at depths $1$ through $d-1$; and $\beta(\cdot)$ is a reshaping function, such as the Benjamini-Yekutieli [44] function $\beta_{BY}(x)=x/\sum_{k=1}^{V}1/k$, which is designed to ensure FDR control under arbitrary dependence.
@@
Appendix C Computational Complexity of Constructing the Reliability Graph
Constructing the RG involves three main steps: training the BT model, clustering, and Lasso regression. The training of the BT model has a time complexity $\mathcal{O}(n^{2})$, because it involves a cyclic optimization over the model parameters associated to the pairs of configurations [13, Sec. 3]. Hierarchical clustering has a complexity $\mathcal{O}(n^{2})$ [45, Sec. 3], while Lasso regression has a per-iteration complexity $\mathcal{O}(n)$ [46, Sec. 2.1]. Therefore, the overall complexity of RG construction is of the order $\mathcal{O}(n^{2})$.
Appendix D RG-PT Algorithm
Algorithm 2 provides a summary of RG-PT.
Algorithm 2 Reliability Graph-Based Pareto Testing (RG-PT)
1:  Input: Hyperparameter set $\Lambda$, calibration dataset $\mathcal{Z}$, FDR level $\delta$, reliability thresholds $\{\alpha_{l}\}_{l=1}^{L_{c}}$, number of DAG levels $D$
2:  Output: Reliable hyperparameter subset $\hat{\Lambda}_{\mathcal{Z}}$
3:  Split calibration data: $\mathcal{Z}=\mathcal{Z}_{\text{OPT}}\cup\mathcal{Z}_{\text{MHT}}$
4:  Estimate Pareto front $\Lambda_{\text{OPT}}\subseteq\Lambda$ using $\mathcal{Z}_{\text{OPT}}$
5:  Construct Reliability Graph (RG):
6:  Compute pairwise comparisons using $\mathcal{Z}_{\text{OPT}}$ and optional priors
7:  Estimate scores $s(\lambda)$ via Bradley-Terry model
8:  Cluster $\Lambda_{\text{OPT}}$ into $D$ levels using scores
9:  for each level $d=2$ to $D$ do
10:     for each $\lambda\in\Lambda_{d}$ do
11:        Select parents in $\Lambda_{d-1}$ via non-negative Lasso
12:     end for
13:  end for
14:  Run DAGGER on RG with data $\mathcal{Z}_{\text{MHT}}$:
15:  for each node $\lambda$ in topological order do
16:     if all parents of $\lambda$ are reliable then
17:        Compute p-value $p_{\lambda}$
18:        if $p_{\lambda}\leq\delta_{\lambda}$ then
19:           Add $\lambda$ to $\hat{\Lambda}_{\mathcal{Z}}$
20:        end if
21:     end if
22:  end for
23:  return  $\hat{\Lambda}_{\mathcal{Z}}$
Appendix E Additional Experiments
In this section, we present four new experiments across language model calibration [27], object detection [4], image classification [47], and telecommunications engineering [28], to demonstrate the advantages of our method over LTT and PT. We begin by detailing the hardware used for the simulations and the specific parameter settings chosen for each experiment.
E.1 Experimental Setups
All experiments were conducted using dedicated computational resources. Specifically, RG-PT, LTT, and PT runs, along with data generation for the object detection, image classification, and telecommunications engineering tasks, were executed on a machine equipped with an Apple M1 Pro chip (10-core CPU, 16-core GPU, 16 GB RAM). Data generation for the prompt engineering experiment (Section 4.1) and the sequence-to-sequence translation task was performed on an NVIDIA A100 GPU (40 GB VRAM), using CUDA 11.3 and 40 GB system memory.
The RG-PT parameter settings for each experiment are summarized in Table 1.
Table 1: RG-PT parameter settings for each experiment.
Experiment $D$ $n_{p}$ $\tau$
Prompt Engineering 17 1000 0.1
Sequence-to-sequence translation 10 200 0.1
Object Detection 20 0 0.1
Image Classification - Low Dimension 10 0 0.1
Image Classification - High Dimension 20 0 0.1
Telecommunications Engineering 10 0 0.1
E.2 FDR Analysis Across All Experiments
To begin with, we present a high-level summary of our experimental validation of Proposition 3.1. The specific details of each experiment, including their objectives, datasets, and task configurations, are provided in their respective sections. Briefly, for each task, we ran RG-PT 100 times using different random splits of the available dataset into $\mathcal{Z}_{\text{OPT}}$, $\mathcal{Z}_{\text{MHT}}$, and an independent test set. The target was to control the average FDR on the test set below a threshold of $\delta=0.1$. We then measured the average FDR on the test set across runs, and the results, summarized in Table 2, confirm that RG-PT satisfies the FDR condition as theoretically established in Proposition 3.1.
Table 2: Average FDR achieved by RG-PT across tasks for a target FDR threshold of $\delta=0.1$.
Task Prompt Engineering Object Detection Language Translation Image Classification Radio Access Scheduling
Average FDR 0.089 0.093 0.095 0.084 0.093
E.3 Ablation Study
In this section, we use the prompt engineering task in Sec. 4.1 to perform an ablation study over the RG depth $D$, as well as the effect of misspecified prior information.
E.3.1 Effect of DAG Depth
To assess the impact of the DAG depth $D$ in RG-PT, we vary it from 1 (flat graph, equivalent to LTT) to 100 (fully linear, equivalent to PT), and measure both the length of the shortest prompt in $|\hat{\Lambda}_{\mathcal{Z}}|$ and the test FDR. Fig. 6 illustrates the average shortest reliable prompt length and average FDR over 100 runs for each depth, showing that FDR remains valid across depths. Additionally, it can be seen that there exists an intermediate depth that minimizes the average prompt length, indicating that the depth $1<D<|\Lambda_{\text{OPT}}|$ can be chosen to optimize power.
Figure 6: Average shortest prompt length in the returned prompt set $\hat{\Lambda}_{\mathcal{Z}}$ and the average FDR achieved by RG-PT as a function of RG depth $D$.
E.3.2 Effect of Misspecified Priors
Figure 7: Average shortest prompt length in the returned prompt set $\hat{\Lambda}_{\mathcal{Z}}$ and the average FDR achieved by RG-PT as a function of fraction $f$ of flipped pairwise prior probabilities.
To simulate prior misspecification, we inject noise into the pairwise priors $\eta_{ij}$. For every pair $\lambda_{i},\lambda_{j}\in\Lambda_{\text{OPT}}$, we independently swap the priors $\eta_{ij}$ and $\eta_{ji}$ with probability $f\in[0,1]$. When $f=0$ the prior remains intact, and when $f=1$ every pairwise preference is completely reversed.
Fig. 7 illustrates the average shortest prompt length in the returned set $\hat{\Lambda}_{\mathcal{Z}}$ of reliable prompts and the average FDR, across 100 random data splits, as a function of the flipped fraction $f$ of pairwise preferences $\eta_{ij}$. As $f$ increases, the shortest prompt length grows, indicating the effect of reduced prior information quality. Despite this, the FDR remains controlled below the target level $\delta=0.1$, demonstrating robustness to misspecified priors. Notably, for $f>0.6$, RG-PT returns an empty set $\hat{\Lambda}_{\mathcal{Z}}$ across all runs, correctly avoiding any potentially unreliable hyperparameters in the face of highly corrupted priors.
E.4 Image Segmentation for Object Detection
We now evaluate the proposed RG-PT framework on a multi-objective image segmentation task for object detection, leveraging the MS-COCO dataset [48] and a pretrained detector from Detectron2 [49] as done in [4]. The task involves three distinct objectives: (i) detecting objects within an image (object detection); (ii) delineating object boundaries (image segmentation); and (iii) assigning correct labels to detected objects (object classification). These tasks are measured using recall, intersection-over-union (IoU), and classification accuracy, respectively. The goal is to control classification errors while optimizing recall and segmentation quality, addressing the trade-offs among these objectives.
The candidate hyperparameter set $\Lambda$ was constructed as per [4], by taking all combinations of 50 linearly spaced values in $[0.2,0.5]$ for $\lambda_{1}$, 5 linearly spaced values in $[0.3,0.7]$ for $\lambda_{2}$, and 25 logarithmically spaced values in $[-0.00436,0]$ for $\lambda_{3}$. These discretization choices were optimized [4].
Figure 8: Illustration of the benefits of the proposed RG-PT hyperparameter selection scheme over the state-of-the-art LTT and PT for an object detection application [4]. The red arrows mark the objects not detected by an object recognition model calibrated using LTT or PT that are instead detected by the same model calibrated via RG-PT (see Appendix E.4 for details).
Denote as $R_{1}(\lambda)$, $R_{2}(\lambda)$, and $R_{3}(\lambda)$ the risks associated with recall, IOU, and coverage, respectively, for hyperparameter $\lambda=(\lambda_{1},\lambda_{2},\lambda_{3})$. Controlling these risks in the context of problem (3) is equivalent to having $L=3$ and $L_{c}=3$. Additionally, as in [4], we set the targets as $\alpha_{1}=0.5$, $\alpha_{2}=0.5$, and $\alpha_{3}=0.75$ with $\delta=0.1$. Within the set $\hat{\Lambda}_{\mathcal{Z}}$ of reliable hyperparameters returned by the algorithm of choice, we choose the hyperparameter in subset $\hat{\Lambda}_{\mathcal{Z}}$ with the lowest value of $\lambda_{1}$ in order to increase the number of detected objects as much as possible. No prior knowledge was leveraged in creating the RG in this experiment by setting $n_{p}=0$.
E.5 Image Classification
@@
In this experiment, following [47], we consider the problem of hyperparameter selection for a support vector machine (SVM) model used to classify images from the Fashion MNIST dataset. The Fashion MNIST is a widely used benchmark for image classification, consisting of 70,000 grayscale images of 10 different clothing categories.
@@
This is a non-trivial problem since the accuracy maximizing model may not also optimize the recall [50].
The SVM model requires the selection of two hyperparameters [47]. The regularization parameter, $C$, controls the desired trade-off between maximizing the margin and minimizing the classification error. Lower values of $C$ allow for a softer margin that can overlook some misclassification errors, while higher values enforce stricter classification error requirements. The kernel coefficient, $\gamma$, determines the impact of a single training example on the decision boundary, with higher values capturing finer details in the data set but risking overfitting. To create the initial candidate set $\Lambda$, we selected hyperparameters over a grid of 25 combinations, using five logarithmically spaced values in the intervals $[-3,3]$ and $[-4,1]$ for $C$ and $\gamma$, respectively. This selection is in line with approaches such as [51] for SVM hyperparameter selection.
@@
Fig. 10 illustrates the testing procedures of LTT [4], PT [5], and RG-PT. The x- and y-axes represent the logarithmic scales of the two hyperparameters $C$ and $\gamma$, while the contours indicate levels of recall $R_{\text{rec}}(\lambda)$, evaluated on the test data set, as a function of the hyperparameters $\lambda=(c,\gamma)$. The numbers illustrate the testing order for each testing method. Note that than LTT, which uses BY, does not follow any inferred order on the hyperparameters, and thus does not have the order labels in the figures. Furthermore, while LTT and PT test hyperparameters one by one, following a linear trajectory, RG-PT proceeds along a DAG, testing at the same time all hyperparameters at the same depth in the DAG.
@@
To demonstrate the scalability of RG-PT to high-dimensional hyperparameter spaces, we repeated the previous experiment over a grid of 10,000 hyperparameter configurations instead of 25. This grid was constructed using 100 logarithmically spaced values for $C$ and $\gamma$ over the same intervals as before. The average FDR across 100 runs is reported in Table 2, highlighting RG-PT’s robustness and effectiveness even in high-dimensional settings.
@@
Figure 10: Illustration of the hyperparameter selection procedure followed by LTT (a), PT (b), and RG-PT (c) for the setting studied in Sec. E.5. Each node represents a hyperparameter $\lambda=(C,\gamma)$, with the numbers representing the testing order. Green nodes show the hyperparameters included in the reliable set $\hat{\Lambda}_{\mathcal{Z}}$, and the star node shows the hyperparameter in set $\hat{\Lambda}_{\mathcal{Z}}$ with the lowest recall rate.
E.6 Radio Access Scheduling
In this section, we study a telecommunications engineering problem, namely the optimization of a radio access scheduler [28]. In this setup, each user equipment (UE) belongs to one of four quality-of-service (QoS) classes, assigned at random, each with its own delay and bit rate requirements [52]. The goal is to control the delay of UEs in a given QoS class, while simultaneously minimizing the delays for UEs in the other three QoS classes.
@@ E.6 Radio Access Scheduling
The scheduling algorithm at the base station allocates spectral resources to the UEs. As in [28], the UEs are randomly distributed within a $1\,\text{km}^{2}$ area containing a centrally located base station. Each UE has an initial buffer of 100 packets, and moves at random speeds and directions. Resource allocation is carried out in intervals of 1 ms, called transmission time intervals (TTIs), over 10,000 TTIs per episode.
The resource allocation algorithm is controlled by a set of hyperparameters $\lambda=(\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4})\in\Lambda$, where each $\lambda_{i}$ adjusts a specific criterion in the reward model as detailed in [28]. Hyperparameters $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$, and $\lambda_{4}$ determine respectively the channel quality for each UE, the total queue sizes at the UEs, the age of the oldest packet in each UE’s buffer, and the fairness in resource block allocation among UEs.
Calibration and test data were generated using the Nokia wireless suite [28]. For each run, we used 100 episodes for calibration and 100 episodes for testing.
The candidate hyperparameter set $\Lambda$ was generated by keeping $\lambda_{1}$ and $\lambda_{2}$ at the values $\lambda_{1}^{*}$ and $\lambda_{2}^{*}$ recommended by [52], and linearly sweeping hyperparameters $\lambda_{3}$ and $\lambda_{4}$ in [0.02, 0.2] and [-0.1, 0.1], respectively, with 10 steps each, resulting in a total of 100 combinations.